15x-10y=-5
15y=4+5x
: 15x-10y=-5
15y=4+5x

Answer by checkley77(3848)

15y=4+5x or -5x+15y=4 now multiply this equation by 3 & add them
15x-10y=-5
-15x+45y=12
---------------------
35y=7
y=7/35
y=1/5 ans.
15x+45*1/5=12
15x+9=12
15x=12-9
15x=3
x=3/15
x=1/5 ans.
Proof:
15*1/5=4-5*1/5
15/5=4-5/5
3=4-1
3=3